# Group-Adaptive Threshold Optimization for Robust AI-Generated Text Detection

### v3

@@ Appendix C Applied Thresholds
Table 4: Thresholds that used in the study
Appendix D Feature Engineering
@@
In order to mitigate disparities within AI classifiers, a thorough feature engineering process was undertaken. Features were systematically extracted to characterize each text within the dataset. The table 5 includes all extraced features for this research. The primary focus of the experiment encompassed text length and anticipated personality traits. Although formality and sentiment were also extracted, these attributes are utilized to expedite processing, as the computational time increases exponentially with the inclusion of additional features. The subsequent table illustrates the features primarily extracted for the experiment, with an emphasis solely on the utilization of length and personality traits.
@@
The text length attribute measures the number of characters in a text. OpenAI’s evaluations (OpenAI, 2023) revealed performance disparities between short (under 1,000 characters) and long texts when using their AI classifier. While their global threshold approach failed to address group-specific differences, it provided a valuable benchmark. Building on this and using exploratory data analysis (EDA), we refined the thresholds to 1,000 and 2,500 characters, categorizing texts as short (up to 1,000), mid-length (1,000–2,500), or long (over 2,500).
@@
The personality attribute includes five traits: extroversion, neuroticism, agreeableness, conscientiousness, and openness. This attribute contributes to the enforcement of fairness metrics, such as demographic parity and equality of odds. For this feature, we determine the personality trait by selecting the one with the highest probability among the five using the (Minej, 2023) model, which is fine-tuned from BERT-base-uncased (Devlin et al., 2018).
Attribute Characteristics Labels
Text Length Measures the number of characters in a text short, medium, long (n = 3)
Formality Indicates the level of formality in the language formal, informal (n = 2)
Sentiment Reflects the emotional tone. positive, negative (n= 2)
Personality Represents the inferred Big Five personality traits. Openness, Conscientiousness, Extroversion, Agreeableness, Neuroticism (n = 5)
Table 5: Characteristics and labels of each attribute
Outside of the length and sentiment, the formality attribute measures a text’s formality by analyzing the frequency of specific parts of speech. According to (Heylighen & Dewaele, 2002), with formal texts typically containing a higher proportion of nouns, adjectives, articles, and prepositions, whereas informal texts tend to have a greater frequency of pronouns, adverbs, verbs, and interjections. As general rule, we set a threshold of 50 to distinguish between formal (scores above 50) and informal texts.
The sentiment attribute allows for the assessment of emotional tone and attitudes within content. AI-generated texts may display distinct sentiment patterns due to the differences in their underlying algorithms compared to human expression. To classify texts as positive or negative, we used the label output from the (DistilBERT-community, ) model, which is a fine-tuned version of DistilBERT-base-uncased (Sanh et al., 2019), specifically trained on the SST-2 dataset(Socher et al., 2013).
Appendix E Discrepancies by classifiers with additional features
E.1 Discrepancy of Probability in Subgroups
E.1.1 Discrepancy in RoBERTa-large Detector
@@
The analysis revealed significant variability in the probability distributions between subgroup pairs, as depicted in Figure 7. Employing two-sample Kolmogorov-Smirnov (KS) statistic for two-sided tests, we assessed distributional discrepancies across various text groups characterized by distinct feature combinations, under the assumption of a non-parametric distribution. Significant differences were identified in certain subgroup comparisons, indicating notable discrepancies in their probability distributions. For example, comparing texts characterized by conscientiousness, positive sentiment, formal style, and long length with those exhibiting extroversion, negative sentiment, formal style, and long length resulted in a KS statistic of 0.7286 ($p<0.05$). Similarly, evaluating agreeableness, positive sentiment, formal style, and mid-length against conscientiousness, positive sentiment, formal style, and long length yielded a KS statistic of 0.7143 ($p<0.01$).
Conversely, smaller discrepancies were observed between neuroticism, positive sentiment, formal style, and short length versus openness, positive sentiment, formal style, and mid-length (KS = 0.2706, $p<0.05$) and neuroticism, negative sentiment, formal style, and short length versus openness, positive sentiment, formal style, and mid-length (KS = 0.2323, $p<0.05$).
Figure 7: Kernel Density Estimation (KDE) plots comparing probability distributions for subgroup pairs with the top two highest and lowest discrepancies based on the K-S test with statistical validity. The largest discrepancy is observed with a KS statistic of $0.7286$, while the smallest discrepancy has a KS statistic of $0.2323$.
E.1.2 Discrepancy in RoBERTa-base Detector
The analysis revealed a substantial variation in the probability distributions among pairs of subgroups, as shown in Figure 8. The KS test was applied to assess distributional differences across diverse text groupings defined by multiple feature sets, under the assumption of a non-parametric distribution. Specific subgroup comparisons highlighted significant differences, pointing to noticeable disparities in their probability distributions. For example, contrasting texts defined by agreeableness, positive sentiment, formal style, and mid-length with those featuring neuroticism, negative sentiment, informal style, and short length produced a KS statistic of 0.9 ($p<0.01$). Furthermore, the comparison between conscientiousness, negative sentiment, formal style, and long length and neuroticism, negative sentiment, informal style, and short length resulted in a KS statistic of 0.7750 ($p<0.01$).
Figure 8: KDE plots comparing probability distributions for subgroup pairs with the highest and lowest discrepancies based on the K-S test with statistical validity. The largest discrepancy is observed with a KS statistic of $0.9$, while the smallest discrepancy has a KS statistic of $0.5071$. These discrepancies indicate varying levels of divergence between groups based on their characteristics.
Lesser discrepancies were noted between neuroticism, negative sentiment, informal style, and short length versus openness, positive sentiment, formal style, and long length (KS = 0.5522, $p<0.05$) and neuroticism, negative sentiment, informal style, and short length versus openness, positive sentiment, formal style, and short length (KS = 0.5071, $p<0.05$).
E.1.3 Discrepancy in RADAR Detector
The investigation disclosed considerable variability in the probability distributions across subgroup pairs, as illustrated in Figure 9. Through the application of the KS test, we evaluated distributional discrepancies among various text groups defined by distinct feature combinations, operating under the premise of a non-parametric distribution. Significant differences were discerned in particular subgroup comparisons, which highlight notable discrepancies in their probability distributions. For instance, the comparison between texts characterized by conscientiousness, negative sentiment, informal style, and long length and those exhibiting extroversion, positive sentiment, formal style, and short length yielded a KS statistic of 0.8 ($p<0.05$).
Figure 9: KDE plots comparing probability distributions for subgroup pairs with the highest and lowest discrepancies based on the K-S test with statistical validity. The largest discrepancy is observed with a KS statistic of $0.8$, while the smallest discrepancy has a KS statistic of $0.1830$.
Similarly, the examination of conscientiousness, negative sentiment, informal style, and long length against neuroticism, negative sentiment, formal style, and short length resulted in a KS statistic of 0.6667 ($p<0.05$). In contrast, smaller discrepancies were detected between conscientiousness, negative sentiment, formal style, and long length versus openness, positive sentiment, formal style, and mid-length (KS = 0.2110, $p<0.01$) and openness, negative sentiment, formal style, and long length versus openness, positive sentiment, formal style, and mid-length (KS = 0.1830, $p<0.01$).
E.1.4 Discrepancy in GPT4o-mini Detector
The analysis revealed substantial variability in the probability distributions among subgroup pairs, as depicted in Figure 10. Using the KS test, we assessed distributional discrepancies across various text groups defined by distinct feature combinations, assuming a non-parametric distribution. Significant differences were detected in particular subgroup comparisons, suggesting notable divergences in their probability distributions. For example, the comparison between texts characterized by openness, negative sentiment, informal style, and mid-length and those exhibiting openness, positive sentiment, informal style, and mid-length resulted in a KS statistic of 0.6667 ($p<0.001$). Similarly, the evaluation of extroversion, negative sentiment, formal style, and mid-length versus openness, positive sentiment, informal style, and mid-length produced a KS statistic of 0.6522 ($p<0.001$). Conversely, smaller discrepancies were noted between neuroticism, positive sentiment, formal style, and short length versus openness, negative sentiment, formal style, and long length (KS = 0.2853, $p<0.05$) and neuroticism, negative sentiment, formal style, and short length versus openness, negative sentiment, formal style, and long length (KS = 0.2560, $p<0.05$).
Figure 10: KDE plots comparing probability distributions for subgroup pairs with the highest and lowest discrepancies based on the K-S test with statistical validity. The largest discrepancy is observed with a KS statistic of $0.6667$, while the smallest discrepancy has a KS statistic of $0.2560$.
Appendix F Operationalizing FairOPT
F.1 Algorithmic description

### v4

@@ Appendix D Applied Thresholds
Table 3: Thresholds from three thresholding methods that used in the experiment.
Appendix E Detailed Results
We report a table showing the overall discrepancy, ACC, and F1 by dataset source: Table 4, and a table showing the detailed difference by detectors: Table LABEL:tab:overall_by_detector_and_dataset.
E.1 Overall discrepancy and performance
The table 4 summarizes the overall discrepancy and performance by each dataset and thresholding method. Detailed results that include the result from each detector is noted in the next section.
Source Strategy Discrepancy (mean $\pm$ std) Accuracy (mean $\pm$ std) F1 Score (mean $\pm$ std)
MAGE FairOPT $\mathbf{0.4495\pm 0.1656}$ $0.6339\pm 0.1512$ $0.3877\pm 0.3494$
ROCFPR $0.5551\pm 0.1664$ $0.6238\pm 0.2113$ $0.5326\pm 0.3170$
Static $0.5433\pm 0.1781$ $0.6388\pm 0.2015$ $0.5126\pm 0.3352$
RAID FairOPT $\mathbf{0.4836\pm 0.1781}$ $0.6280\pm 0.1573$ $0.3885\pm 0.3535$
ROCFPR $0.5644\pm 0.1263$ $0.6253\pm 0.2082$ $0.5363\pm 0.3185$
Static $0.5562\pm 0.1478$ $0.6364\pm 0.1994$ $0.5136\pm 0.3358$
SemEval FairOPT $\mathbf{0.3763\pm 0.2166}$ $0.6404\pm 0.1497$ $0.3896\pm 0.3484$
ROCFPR $0.5598\pm 0.1758$ $0.6258\pm 0.2010$ $0.5303\pm 0.3114$
Static $0.5548\pm 0.1969$ $0.6376\pm 0.1948$ $0.5081\pm 0.3288$
Table 4: Overall discrepancy and performance by data source and thresholding strategy (mean $\pm$ std)
E.2 Comparison by performance and fairness metrics by datasource, detector, and thresholding strategy
The table LABEL:tab:overall_by_detector_and_dataset shows the results by each detector by the datasource and thresholding method based on fairness and performance metrics.
Table 5: Discrepancy and performance (mean $\pm$ std) by data source, detector, and thresholding method
Dataset
Detector Thresholding $Discrepancy$ $ACC$ $F1score$
MAGE
GPT4.1-nano (2025) FairOPT $0.4106\pm 0.2350$ $0.4794\pm 0.0374$ $0.0114\pm 0.0121$
ROCFPR $0.4636\pm 0.2194$ $0.4274\pm 0.0380$ $0.0694\pm 0.0159$
Static $0.4647\pm 0.2143$ $0.4393\pm 0.0401$ $0.0507\pm 0.0154$
GPT4o-mini (2024) FairOPT $0.3820\pm 0.2421$ $0.4866\pm 0.0345$ $0.0046\pm 0.0078$
ROCFPR $0.6874\pm 0.1882$ $0.4453\pm 0.0288$ $0.2348\pm 0.0399$
Static $0.6517\pm 0.1661$ $0.4423\pm 0.0253$ $0.1947\pm 0.0304$
SimpleAI (2023) FairOPT $0.4536\pm 0.0871$ $0.7501\pm 0.0274$ $0.7006\pm 0.0363$
ROCFPR $0.5682\pm 0.0945$ $0.7573\pm 0.0130$ $0.7520\pm 0.0215$
Static $0.5527\pm 0.0968$ $0.7525\pm 0.0154$ $0.7360\pm 0.0298$
DeskLib (2024) FairOPT $0.4896\pm 0.0914$ $0.8900\pm 0.0170$ $0.8809\pm 0.0196$
ROCFPR $0.4854\pm 0.2388$ $0.8787\pm 0.0212$ $0.8916\pm 0.0224$
Static $0.4278\pm 0.2561$ $0.9074\pm 0.0206$ $0.9140\pm 0.0219$
E5LoRA (2024) FairOPT $0.3571\pm 0.2440$ $0.4895\pm 0.0336$ $00.0000\pm 0.0$ 0.0000
ROCFPR $0.5779\pm 0.1866$ $0.1996\pm 0.0302$ $0.0471\pm 0.0179$
Static $0.6026\pm 0.2439$ $0.2797\pm 0.0237$ $0.0065\pm 0.0090$
FastDetectGPT (2023) FairOPT $0.4929\pm 0.0189$ $0.8099\pm 0.0274$ $0.7774\pm 0.0243$
ROCFPR $0.4881\pm 0.0315$ $0.7776\pm 0.0302$ $0.8114\pm 0.0288$
Static $0.4911\pm 0.1295$ $0.8398\pm 0.0154$ $0.8520\pm 0.0151$
RoBERTa-base (2019) FairOPT $0.4893\pm 0.0537$ $0.6742\pm 0.0436$ $0.5381\pm 0.0655$
ROCFPR $0.5923\pm 0.0884$ $0.7322\pm 0.0311$ $0.6963\pm 0.0320$
Static $0.6158\pm 0.1606$ $0.7256\pm 0.0364$ $0.6722\pm 0.0423$
RoBERTa-large (2019) FairOPT $0.3940\pm 0.1992$ $0.5063\pm 0.0364$ $0.0656\pm 0.0372$
ROCFPR $0.5835\pm 0.0610$ $0.6264\pm 0.0289$ $0.5067\pm 0.0405$
Static $0.5414\pm 0.1038$ $0.6103\pm 0.0361$ $0.4315\pm 0.0522$
RADAR (2023) FairOPT $0.5760\pm 0.0755$ $0.6192\pm 0.0216$ $0.5106\pm 0.0388$
ROCFPR $0.5495\pm 0.2113$ $0.7693\pm 0.0186$ $0.7837\pm 0.0202$
Static $0.5418\pm 0.1280$ $0.7519\pm 0.0224$ $0.7562\pm 0.0242$
RAID
GPT4.1-nano (2025) FairOPT $0.5208\pm 0.2082$ $0.4738\pm 0.0087$ $0.0128\pm 0.0115$
ROCFPR $0.5820\pm 0.1623$ $0.4243\pm 0.0189$ $0.0795\pm 0.0295$
Static $0.5781\pm 0.1558$ $0.4306\pm 0.0183$ $0.0465\pm 0.0202$
GPT4o-mini (2024) FairOPT $0.4781\pm 0.2428$ $0.4710\pm 0.0180$ $00.0000\pm 0.0$ 0.0000
ROCFPR $0.5762\pm 0.1567$ $0.4557\pm 0.0266$ $0.2292\pm 0.0542$
Static $0.5816\pm 0.1733$ $0.4529\pm 0.0180$ $0.1875\pm 0.0321$
SimpleAI (2023) FairOPT $0.5868\pm 0.2080$ $0.7320\pm 0.0408$ $0.6835\pm 0.0548$
ROCFPR $0.5655\pm 0.1818$ $0.7452\pm 0.0386$ $0.7502\pm 0.0388$
Static $0.6245\pm 0.2095$ $0.7403\pm 0.0516$ $0.7357\pm 0.0541$
DeskLib (2024) FairOPT $0.3552\pm 0.1001$ $0.9023\pm 0.0289$ $0.8973\pm 0.0301$
ROCFPR $0.4578\pm 0.0522$ $0.8667\pm 0.0223$ $0.8843\pm 0.0194$
Static $0.4292\pm 0.0646$ $0.8988\pm 0.0212$ $0.9076\pm 0.0189$
E5LoRA (2024) FairOPT $0.4167\pm 0.2041$ $0.4808\pm 0.0153$ $00.0000\pm 0.0$ 0.0000
ROCFPR $0.6437\pm 0.1177$ $0.2059\pm 0.0384$ $0.0406\pm 0.0272$
Static $0.6535\pm 0.1981$ $0.2805\pm 0.0323$ $0.0078\pm 0.0097$
FastDetectGPT (2023) FairOPT $0.4405\pm 0.0560$ $0.8241\pm 0.0203$ $0.7996\pm 0.0230$
ROCFPR $0.5103\pm 0.0873$ $0.7850\pm 0.0345$ $0.8191\pm 0.0312$
Static $0.4609\pm 0.1409$ $0.8472\pm 0.0247$ $0.8601\pm 0.0228$
RoBERTa-base (2019) FairOPT $0.5361\pm 0.0452$ $0.6706\pm 0.0343$ $0.5483\pm 0.0681$
ROCFPR $0.6246\pm 0.1244$ $0.7348\pm 0.0355$ $0.7030\pm 0.0478$
Static $0.5978\pm 0.0984$ $0.7180\pm 0.0342$ $0.6682\pm 0.0499$
RoBERTa-large (2019) FairOPT $0.4282\pm 0.1927$ $0.4990\pm 0.0183$ $0.0698\pm 0.0285$
ROCFPR $0.6079\pm 0.0816$ $0.6413\pm 0.0218$ $0.5341\pm 0.0394$
Static $0.5570\pm 0.0658$ $0.6176\pm 0.0175$ $0.4579\pm 0.0319$
RADAR (2023) FairOPT $0.5900\pm 0.1850$ $0.5980\pm 0.0179$ $0.4853\pm 0.0412$
ROCFPR $0.5113\pm 0.0602$ $0.7683\pm 0.0263$ $0.7867\pm 0.0233$
Static $0.5232\pm 0.0644$ $0.7418\pm 0.0265$ $0.7510\pm 0.0257$
SemEval
GPT4.1-nano (2025) FairOPT $0.2663\pm 0.2792$ $0.4889\pm 0.0498$ $0.0110\pm 0.0116$
ROCFPR $0.5139\pm 0.2100$ $0.4459\pm 0.0354$ $0.0912\pm 0.0320$
Static $0.4889\pm 0.2156$ $0.4495\pm 0.0408$ $0.0503\pm 0.0210$
GPT4o-mini (2024) FairOPT $0.2763\pm 0.2923$ $0.4943\pm 0.0501$ $0.0029\pm 0.0076$
ROCFPR $0.5777\pm 0.2485$ $0.4638\pm 0.0350$ $0.2381\pm 0.0484$
Static $0.5617\pm 0.2843$ $0.4620\pm 0.0418$ $0.2008\pm 0.0364$
SimpleAI (2023) FairOPT $0.5655\pm 0.2123$ $0.7382\pm 0.0369$ $0.6842\pm 0.0260$
ROCFPR $0.6726\pm 0.1674$ $0.7430\pm 0.0313$ $0.7375\pm 0.0264$
Static $0.6519\pm 0.1904$ $0.7436\pm 0.0327$ $0.7283\pm 0.0231$
DeskLib (2024) FairOPT $0.3711\pm 0.0755$ $0.8942\pm 0.0349$ $0.8844\pm 0.0316$
ROCFPR $0.4487\pm 0.0727$ $0.8643\pm 0.0169$ $0.8776\pm 0.0241$
Static $0.3691\pm 0.0970$ $0.8954\pm 0.0139$ $0.9012\pm 0.0184$
E5LoRA (2024) FairOPT $0.2143\pm 0.2673$ $0.5015\pm 0.0523$ $00.0000\pm 0.0$ 0.0000
ROCFPR $0.6887\pm 0.2807$ $0.2098\pm 0.0405$ $0.0293\pm 0.0124$
Static $0.6428\pm 0.2790$ $0.2750\pm 0.0402$ $0.0065\pm 0.0061$
FastDetectGPT (2023) FairOPT $0.4818\pm 0.0312$ $0.8111\pm 0.0254$ $0.7701\pm 0.0298$
ROCFPR $0.4689\pm 0.0646$ $0.7490\pm 0.0284$ $0.7851\pm 0.0345$
Static $0.5232\pm 0.1360$ $0.8159\pm 0.0253$ $0.8268\pm 0.0276$
RoBERTa-base (2019) FairOPT $0.4922\pm 0.0426$ $0.6940\pm 0.0371$ $0.5676\pm 0.0472$
ROCFPR $0.5748\pm 0.1175$ $0.7519\pm 0.0241$ $0.7127\pm 0.0362$
Static $0.5543\pm 0.0866$ $0.7412\pm 0.0263$ $0.6847\pm 0.0435$
RoBERTa-large (2019) FairOPT $0.2956\pm 0.2093$ $0.5188\pm 0.0513$ $0.0718\pm 0.0086$
ROCFPR $0.5977\pm 0.1239$ $0.6479\pm 0.0218$ $0.5280\pm 0.0376$
Static $0.6073\pm 0.1346$ $0.6186\pm 0.0293$ $0.4333\pm 0.0302$
RADAR (2023) FairOPT $0.4241\pm 0.1244$ $0.6222\pm 0.0358$ $0.5140\pm 0.0407$
ROCFPR $0.4956\pm 0.0187$ $0.7561\pm 0.0310$ $0.7735\pm 0.0357$
Static $0.5939\pm 0.1847$ $0.7370\pm 0.0300$ $0.7407\pm 0.0361$
E.3 Results for Discrepancy, ACC, and F1
The bar charts illustrate the discrepancy, ACC, and F1 are incorporated in the main body of the paper, and this section encompasses a chart representing the F1 score, accompanied by three figures. The sequence of figures is arranged in the order of discrepancy, ACC, and F1.
Figure 10: FairOPT decreases the discrepancy, Balanced Error Rate (BER) discrepancy, of probabilistic classifiers compared to other thresholding methods. FairOPT reduces classification discrepancy to $\mu=0.4341\pm 0.1927$, outperforming ROCFPR ($\mu=0.5595\pm 0.1582$) and Static ($\mu=0.5512\pm 0.1758$).
Figure 11: FairOPT shows a negligible sacrifice of ACC by effectively handling the performance-fairness tradeoff using a relaxed fairness that is shown in Figure 4. Our method incurs only a negligible accuracy sacrifice, achieving $\mu=0.6344\pm 0.1518$ compared to Static $\mu=0.6376\pm 0.1975$ and ROCFPR $\mu=0.6249\pm 0.2057$.
Figure 12: Overall F1 have been sacrificed in the application of FairOPT. FairOPT achieves $\mu=0.389\pm 0.350$, compared to ROCFPR $\mu=0.533\pm 0.316$ and Static $\mu=0.511\pm 0.333$. Detectors with low ACC show a very low F1 score.
Appendix F Feature Engineering
@@
In order to mitigate disparities within AI classifiers, a thorough feature engineering process was undertaken. Features were systematically extracted to characterize each text within the dataset. The primary focus of the experiment encompassed text length and anticipated personality traits. Although formality and sentiment were also extracted, these attributes are utilized to expedite processing, as the computational time increases exponentially with the inclusion of additional features. The subsequent table illustrates the features primarily extracted for the experiment, with an emphasis solely on the utilization of length and personality traits.
@@
The text length attribute measures the number of characters in a text. OpenAI’s evaluations [26] revealed performance disparities between short (under 1,000 characters) and long texts when using their AI classifier. While their global threshold approach failed to address group-specific differences, it provided a valuable benchmark. Building on this and using exploratory data analysis (EDA), we refined the thresholds to 1,000 and 2,500 characters, categorizing texts as short (up to 1,000), mid-length (1,000–2,500), or long (over 2,500).
@@
The personality attribute includes five traits: extroversion, neuroticism, agreeableness, conscientiousness, and openness. This attribute contributes to the enforcement of fairness metrics, such as demographic parity and equality of odds. For this feature, we determine the personality trait by selecting the one with the highest probability among the five using the [45] model, which is fine-tuned from BERT-base-uncased [47].
Appendix G Discrepancies by classifiers with additional features
G.1 Discrepancy of Probability in Subgroups
G.1.1 Discrepancy in RoBERTa-large Detector
@@
The analysis revealed significant variability in the probability distributions between subgroup pairs, as depicted in Figure 13. Employing two-sample Kolmogorov-Smirnov (KS) statistic for two-sided tests, we assessed distributional discrepancies across various text groups characterized by distinct feature combinations, under the assumption of a non-parametric distribution. Significant differences were identified in certain subgroup comparisons, indicating notable discrepancies in their probability distributions. For example, comparing texts characterized by conscientiousness and short with those exhibiting neuroticism and long resulted in a KS statistic of 0.3081 ($p<0.01$). Similarly, evaluating conscientiousness and short against extroversion and medium yielded a KS statistic of 0.2616 ($p<0.01$). Conversely, smaller discrepancies were observed between agreeableness and short versus neuroticism and short (KS = 0.0909, $p<0.05$) and neuroticism and short versus openness and short (KS = 0.0842, $p<0.01$).
Figure 13: Top two highest and lowest distributional differences using Kolmogorov-Smirnoff (KS) test on RoB33ERTa-large detector results in cumulative probability distribution function (CDF) based on the training dataset in the range of 0.1 to 0.9. The biggest discrepancy is observed with a KS statistic of $0.3081$ ($p<0.01$), while the smallest is $0.0842$ ($p<0.01$). Statistically significant divergences between specific groups call attention to the need for thresholding methods that are accountable to the varying distributions
Appendix H Operationalizing FairOPT
H.1 Algorithmic description